OpenStudy (anonymous):

Jacob kicks a soccer ball off the ground and in the air, with an initial velocity of 33 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.1 feet 16.5 feet 17.0 feet 18.2 feet

OpenStudy (tylerd):

the maximum height = the vertex Y value

OpenStudy (anonymous):

is the ANSWER C?

OpenStudy (tylerd):

we cant just give away answers and id have to check first

OpenStudy (anonymous):

CAN U CHECK PLS

OpenStudy (tylerd):

im assuming the v=33ft/s

OpenStudy (tylerd):

not sure what the s means

OpenStudy (anonymous):

Hey, I have flvs too! are you in shkoler's class?

OpenStudy (anonymous):

WHY

OpenStudy (anonymous):

im not i have a diff teacher

OpenStudy (anonymous):

have u done this before

OpenStudy (tylerd):

are you sure its + s? or + 5

OpenStudy (anonymous):

aw boo, I'm not done yet, which teacher do you have?

OpenStudy (anonymous):

idk i copied and pasted this

OpenStudy (anonymous):

i forgot her name but what module are you on

OpenStudy (anonymous):

you forgot your teacher's name?

OpenStudy (anonymous):

yeah sorta what module u on

OpenStudy (aum):

H(t) = -16t^2 + vt + s "Jacob kicks a soccer ball OFF THE GROUND" implies the initial height of the ball is 0. v = 33 feet per second. h(t) = -16t^2 + 33t ----- (1) Maximum height is reached at the vertex. t at the vertex = -b/2a = -33/(2*(-16)) = 33/32 sec. plug t = 33/32 into (1) and find the max height.

OpenStudy (aum):

Initial height of the ball s = 0.

OpenStudy (tylerd):

ya thats what i figured

OpenStudy (anonymous):

what s the answer

OpenStudy (tylerd):

no

OpenStudy (aum):

plug t = 33/32 into h(t) = -16t^2 + 33t and find h(t)

OpenStudy (anonymous):

i got 17